You're given the data below for 2021 for US economy whose currency is dollar.
Consumption $350 billion. Transfer payments $100 billion. Investment $100 billion. Government purchases $200 billion. Exports $50 billion. Imports $150 billion. Bond purchases $200 billion. Earnings on foreign investments $75 billion. Foreign earnings on US investment $25 billion.
Compute net foreign investment. Compute net exports. Compute GDP. Compute GNP.
In addition to responding with a quantitative answer, briefly describe how you arrive at your answers.
1
Expert's answer
2021-07-27T02:47:04-0400

Net foreign investment NFI = 75 - 25 = $50 billion.

Net exports NX = 50 - 150 = -$100 billion.

GDP = 350 + 100 + 200 - 100 = $550 billion.

GNP = GDP + NFI = 550 + 50 = $600 billion.